OpenStudy (anonymous):

suppose a superball dropped from a height of 45 ft always rebounds 1/5 of the distance fallen. Approximately how far does it travel before coming to rest? The answer is 67.5, but I have no idea how to get that answer.

OpenStudy (anonymous):

x0 = 1/5*(45) = 9 x1 = 1/5*(9) = 1.8 x2 = 1/5*(1.8) = 0.36 x3 = 0.36/5 = 0.072 45+x0+x1+x2+x3

OpenStudy (surry99):

this is actually just the infinite sum of a geometric series

OpenStudy (mathstudent55):